. a force of 100N is applied to the rim of the pulley wheel of diameter 200mm. Solve for the torque.



1
Expert's answer
2022-01-07T09:29:59-0500

"M=Fl=F\frac d2,"

"M=100\cdot\frac{0.2}2=10~~\text N\cdot\text m."